1.     A student of mass 50kg climbs the college stairs using 2500J of energy, calculate the height of the stairs.


1.     An astronaut of mass 70 kg climbs to a height of 5 m on the moon and gains 560 J of gravitational potential energy. What must be the gravitational field strength on the moon?


1.     What is the speed of a stone of mass 2 kg if it has 36 J of kinetic energy?


1
Expert's answer
2021-05-17T13:31:53-0400

1)

"h=\frac{E}{mg}=5~m,"

2)

"g=\frac{E}{mh}=1.6~\frac{N}{kg},"

3)

"v=\sqrt{\frac{2E}{m}}=6~\frac ms."